OpenStudy (anonymous):

Rewrite as a single logaithm ln3+lny-2lnx

OpenStudy (anonymous):

ln(a) + ln(b) = ln(a*b) and, ln(a) - log(b) = log(a/b) Putting these formulae: ln(3) + ln(y) = ln(3y) Remember, aln(b) = ln(b)^a So, 2ln(x )= ln(x)^2 So, ln(3y) - ln((x)^2 = \[\ln(3y/x^2)\] is the answer...

OpenStudy (anonymous):

is that x^2?

OpenStudy (anonymous):

Yeah, it is x square...

OpenStudy (anonymous):

thank you!message me?